OpenStudy (anonymous):

An object of mass 3.60 kg undergoes a displacement S= 12.2 m i + 7.9 m j while it is being acted on by a net force F = 9.0 N i -1.5 N j. How much work is done on the object by the net force? Also If the object above starts from rest, what is its speed after it has undergone the displacement described above?

OpenStudy (anonymous):

use W = F.S

OpenStudy (anonymous):

dont forget to use dot vector identity

OpenStudy (anonymous):

W = F.s W = (9.0 Ni - 1.5 Nj) . (12.2 mi + 7.9 mj) W = ((9.0 x 12.2)i.i - (1.5x7.9)j.j W = (109.8 - 11.85) N.m

OpenStudy (anonymous):

thank you but how would i put this in loncapa

OpenStudy (anonymous):

Then, F = m.a F = m.a a = F/m \[a =\frac{(9.0 Ni - 1.5 Nj) }{3.60 }\] a = 2.5 N/m i - 0.42 N/m j subtitutes a to vt^2 = vo^2 + 2 a. s, where vt is the final speed, vo is the initial speed, a is acceleration, vt^2 = vo^2 + 2 a. s, vt^2 = 0 + 2 a. s, vt^2 = (2.5 N/m i - 0.42 N/m j) . (12.2 m i + 7.9 m j )

OpenStudy (anonymous):

but i'm sorry I mistakenly wrote: for a unit not N/m but N/kg

OpenStudy (anonymous):

Couldn't you also use kinetic energy to find the speed? That seems more direct than using F=ma and kinematics.

OpenStudy (anonymous):

vt^2 = 2 .(2.5 N/kg i - 0.42 N/kg j) . (12.2 m i + 7.9 m j ) vt^2 = 2 .(30.5 - 3.318) vt = sqrt (2.(30.5 - 3.318)) m/s

OpenStudy (anonymous):

yes i could...,

OpenStudy (anonymous):

W = delta EK

OpenStudy (anonymous):

W = 1/2. mv^2

OpenStudy (anonymous):

Makes the algebra a little easier, eh?

OpenStudy (anonymous):

sure